# Many-Shot In-Context Learning

### v2

@@ 2.3 Planning: Logistics Domain
Figure 5: In-context Planning. Success rate quickly improves with up to 10 shots (37K tokens), followed by saturation up to 400 shots and a sudden performance jump at 800 shots. As a reference, we report 1-shot GPT-4 results from Valmeekam et al. [59]. See Figure A.3 for an example 1-shot prompt.
@@ 4.2 Learning Non-Natural Language Tasks: High-Dimensional Functions
Binary Linear Classification in High Dimensions
Following the setup from Wei et al. [63], we create datasets with $N$-dimensional inputs vectors and their binary class labels, where each dimension is a random integer in $[1,1000]$. See more details in §A.5. While Wei et al. [63] used only 16 shots per class, we scale ICL up to 2048 shots per class. As shown in Figure 11, while $2048$ shots per class perform best when $N=16$, we observe slight accuracy decrease beyond $512$ shots for higher values of $N$ (Figure 11 C, R). Moreover, many-shot ICL substantially outperforms random-chance accuracy and nearly matches the accuracy of a strong baseline, namely $k$-nearest neighbors, indicating that many-shot ICL can implement nearest-neighbour search over inputs. This is reminiscent of induction heads that implement prefix matching over sequences [46], a plausible mechanism for ICL abilities.
Sequential Parity
Parity is a fundamental Boolean function that determines if a binary input sequence contains an even or odd number of 1s. Despite their power, transformers trained specifically for in-context learning, struggle to learn the Parity function over 20-digit sequences [7]. In this work, we evaluate how well many-shot ICL performs with a pretrained LLM to learn the sequential parity function $f(x)=[f_{1}(x),f_{2}(x),\cdots,f_{n}(x)]$, where $x\in\{0,1\}^{n}$ and $f_{i}(x)=x_{1}\oplus x_{2}\cdots\oplus x_{i}\ \forall\ i\in[1,n]$. We report the results in Figure 12. We see consistent improvement in test accuracy as we increase the number of shots to 8192. Performance surpasses a GPT-2 Medium sized transformer [51] trained from scratch on 20$\times$ more input-output examples (with no repeated examples; §A.6). This result indicates many-shot ICL can implement computations analogous to gradient descent [60].
@@
4.3 Is Many-Shot ICL Sensitive to Example Ordering?
In few-shot in-context learning (ICL), the order of examples within the prompt can significantly impact model performance [38, 65]. Here, we investigate whether such sensitivity to prompt ordering observed in few-shot ICL persists in many-shot scenarios, which remains
largely unexplored. Specifically, we evaluate ten different random orderings of fixed 50 in-context examples from MATH training split and evaluate performance on the held-out MATH500 test set [33].
As Figure 13 reveals, performance varies significantly across different subareas in MATH500. Strikingly, an ordering that that excels in one subarea may perform poorly in another, for example, the best Geometry ordering yields weak results on Number Theory. This fluctuation results in a smaller variation in average performance compared to individual subareas. One interesting extension would be to optimize many-shot prompts using frameworks like DSPy [28] that has been successfully applied for optimizing few-shot prompts based a given metric. Overall, these findings highlight a key challenge in ensuring reliable results with many-shot ICL for long-context models.
@@ 4.3 Is Many-Shot ICL Sensitive to Example Ordering?
Many-shot ICL could make task-specific fine-tuning less essential or, in some cases, even unnecessary, allowing LLMs to tackle a wider range of tasks without specialization. While supervised fine-tuning (SFT) is the dominant LLM paradigm when making use of hundreds or thousands of examples, it is computationally expensive in terms of training. In contrast, many-shot ICL does not require any training, however it has a larger inference cost, which can be substantially reduced with KV caching [49, 64], which might be available off-the-shelf with context caching [12].
@@ 6 Discussion, Limitations and Future Work
One limitation of our work is that it mainly evaluates many-shot ICL with Gemini 1.5 Pro. That said, concurrent works [2, 6] as well as our preliminary results with GPT-4-Turbo and Claude-3-Opus (§4.5) indicate that other LLMs can also benefit from many-shot ICL. Future work should focus on evaluating the many-shot abilities of a wide range of long context models, as they become available. Furthermore, many-shot performance can likely serve as a valuable metric for evaluating the quality of long-context models, going beyond the prevalent needle-in-a-haystack test [26].
@@ Acknowledgements
RA initiated and led the project, ran majority of the many-shot experiments and analysis, came up with reinforced ICL, on-boarded collaborators, wrote the initial draft. AS contributed initial infra for experiments on MATH and GSM8K, ran BBH experiments, co-led the fine-tuning experiments, conducted NLL analysis on problem-solving tasks, and wrote several sections.

### v3

@@ 2.3 Planning: Logistics Domain
Figure 5: In-context Planning. A recent version of 1.5 Pro starts from a high few-shot performance, and its many-shot performance scales uniformly from 42% to 62%. For an older version, success rate quickly improves with up to 10 shots (37K tokens), followed by saturation. As a reference, we report 1-shot GPT-4 results [59].
@@ 4.2 Learning Non-Natural Language Tasks: High-Dimensional Functions
Binary Linear Classification in High Dimensions Following the setup from Wei et al. [63], we create datasets with $N$-dimensional inputs vectors and their binary class labels, where each dimension is a random integer in $[1,1000]$. See more details in §A.5. While Wei et al. [63] used only 16 shots per class, we scale ICL up to 2048 shots per class. As shown in Figure 12, while $2048$ shots per class perform best when $N=16$, we observe slight accuracy decrease beyond $512$ shots for higher values of $N$ (Figure 12 C, R). Moreover, many-shot ICL substantially outperforms random-chance accuracy and nearly matches the accuracy of a strong baseline, namely $k$-nearest neighbors, indicating that many-shot ICL can implement nearest-neighbour search over inputs. This is reminiscent of induction heads that implement prefix matching over sequences [46], a plausible mechanism for ICL abilities.
Sequential Parity. Parity is a fundamental Boolean function that determines if a binary input sequence contains an even or odd number of 1s. Despite their power, transformers trained specifically for in-context learning, struggle to learn the Parity function over 20-digit sequences [7]. In this work, we evaluate how well many-shot ICL performs with a pretrained LLM to learn the sequential parity function $f(x)=[f_{1}(x),f_{2}(x),\cdots,f_{n}(x)]$, where $x\in\{0,1\}^{n}$ and $f_{i}(x)=x_{1}\oplus x_{2}\cdots\oplus x_{i}\ \forall\ i\in[1,n]$. We report the results in Figure 12. We see consistent improvement in test accuracy as we increase the number of shots to 8192. Performance surpasses a GPT-2 Medium sized transformer [51] trained from scratch on 20$\times$ more input-output examples (with no repeated examples; §A.6). This result indicates many-shot ICL can implement computations analogous to gradient descent [60].
@@
Many-shot ICL could make task-specific fine-tuning less essential or, in some cases, even unnecessary, allowing LLMs to tackle a wider range of tasks without specialization. While supervised fine-tuning (SFT) is the dominant LLM paradigm when making use of hundreds or thousands of examples, it is computationally expensive in terms of training. In contrast, many-shot ICL does not require any training, however it has a larger inference cost, which can be substantially reduced with KV caching, which might be available off-the-shelf with context caching [12].
@@
4.4 Inference Cost of Many-Shot ICL
Figure 14: Per-Output runtime as we increase shots, averaged across the test set and multiple seeds, on (left) summarization and (right) parity prediction. When computing the next token, we still have to attend to the fixed many-shot prompt, even if KV is cached. We hypothesize that up to a token length of 32K, you can fit the entire KV cache into TPU HBM, which roughly means that you compute next tokens in O(1) memory load.
While many-shot ICL increases inference computation time, it can allow for quick prototyping and experimentation using just an inference API. Moreover, being able to spend additional inference-time compute to obtain better performance is a useful feature to have. With KV caching [49, 64] enabled (default for long-context servers), as shown in Figure 14, runtime increases linearly with a large number of shots, as opposed to quadratically for self-attention: doubling the number of shots nearly doubles the runtime. However, for a small number of shots, runtime is nearly constant. When the number of generated tokens is much smaller than many-shot prompts, each new token is still linear, which explains our observed runtime for a large number of shots.
@@ 4.5 Comparing Many-Shot Abilities of Frontier LLMs
To understand the role of model size, we also evaluated the many-shot performance of Gemini 1.5 Flash, a smaller long-context LLM than Gemini 1.5 Pro. On the English → Bemba task, we find that 1.5 Flash matches Claude-3-Opus and outperforms GPT-4 with 997-shots, despite having much worse few-shot performance than Claude and GPT. On English → Tamil MT, 1.5 Flash outperforms Claude-3 in terms of many-shot performance, while lags behind 1.5 Pro. These results suggest that even smaller LLMs can benefit from many-shot ICL and outperform LLMs with stronger few-shot performance with enough shots.
4.6 Where Do Gains in Many-Shot ICL Stem From?
Figure 16: Many-shot performance with distinct examples vs repeating the same 25 examples $N$ times on low-resource MT. Bars show avg. perf with std across 3 seeds. Most of the benefit of many-shot ICL stems from adding new information.
Including more examples has two effects: (1) increasing information when using distinct samples, (2) but another of increasing the context length. To separate these two effects, we ran an experiment on low-resource MT by repeating 25 examples several times to create many-shot prompts with up to 1000 examples (shuffled ordering) and added the results in Figure 17. The performance with repeated examples stays nearly the same and significantly lags behind many-shot performance with distinct examples. On this task, the benefit of many-shot ICL mainly stems from adding new information as opposed to increasing context length.
4.7 Is Many-Shot ICL Sensitive to Example Ordering?
In few-shot in-context learning (ICL), the order of examples within the prompt can significantly impact model performance [38, 65]. Here, we investigate whether such sensitivity to prompt ordering observed in few-shot ICL persists in many-shot scenarios, which remains largely unexplored. Specifically, we evaluate ten different random orderings of fixed 50 in-context examples from MATH training split and evaluate performance on the held-out 500 examples from MATH test set.
As Figure 17 reveals, performance varies significantly across different splits in MATH. Strikingly, an ordering that that excels in one subarea may perform poorly in another, for example, the best ordering for ‘Split 1’ yields weak results on ‘Split 2’. This fluctuation results in a smaller variation in average performance compared to individual subareas. One interesting extension would be to optimize many-shot prompts using frameworks like DSPy [28] that has been successfully applied for optimizing few-shot prompts based a given metric. Overall, these findings highlight a key challenge in ensuring reliable results with many-shot ICL for long-context models.
@@ 6 Discussion, Limitations and Future Work
One limitation of our work is that it mainly evaluates many-shot ICL with Gemini 1.5 Pro. That said, concurrent works [2, 6] as well as our preliminary results with GPT-4-Turbo, Gemini 1.5 Flash, and Claude-3-Opus (§4.5) indicate that other LLMs can also benefit from many-shot ICL. Future work should focus on evaluating the many-shot abilities of a wide range of long context models, as they become available. Furthermore, many-shot performance can likely serve as a valuable metric for evaluating the quality of long-context models, going beyond the prevalent needle-in-a-haystack test [26].
@@ Acknowledgements
RA initiated and led the project, ran majority of the many-shot experiments and analysis, came up with reinforced ICL, on-boarded collaborators, handled the NeurIPS rebuttal, wrote the initial draft and the revision. AS contributed initial infra for experiments on MATH and GSM8K, ran BBH experiments, co-led the fine-tuning experiments, conducted NLL analysis on problem-solving tasks, and wrote several sections.